40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 15 16:46:27 2021, Anchal, Summary, Optical Levers, Vent Prep Screen_Shot_2021-10-15_at_4.34.45_PM.pngScreen_Shot_2021-10-15_at_4.45.31_PM.pngScreen_Shot_2021-10-15_at_4.34.45_PM.pngScreen_Shot_2021-10-15_at_4.34.45_PM.png
    Reply  Fri Oct 15 17:17:51 2021, Koji, Summary, General, Vent Prep 7x
       Reply  Fri Oct 15 20:53:49 2021, Koji, Summary, General, Vent Prep 
Message ID: 16408     Entry time: Fri Oct 15 17:17:51 2021     In reply to: 16407     Reply to this: 16409
Author: Koji 
Type: Summary 
Category: General 
Subject: Vent Prep 

I took over the vent prep: I'm going through the list in [ELOG 15649] and [ELOG 15651]. I will also look at [ELOG 15652] at the day of venting.

  1. IFO alignment: Two arms are already locking. The dark port beam is well overlapped. We will move PRM/SRM etc. So we don't need to worry about them. [Attachment 1]
    scripts>z read C1:SUS-BS_PIT_BIAS C1:SUS-BS_YAW_BIAS
    -304.7661529521767
    -109.23924626857811
    scripts>z read C1:SUS-ITMX_PIT_BIAS C1:SUS-ITMX_YAW_BIAS
    15.534616817500943
    -503.4536332290159
    scripts>z read C1:SUS-ITMY_PIT_BIAS C1:SUS-ITMY_YAW_BIAS
    653.0100945988496
    -478.16260735781225
    scripts>z read C1:SUS-ETMX_PIT_BIAS C1:SUS-ETMX_YAW_BIAS
    -136.17863332517527
    181.09285307121306
    scripts>z read C1:SUS-ETMY_PIT_BIAS C1:SUS-ETMY_YAW_BIAS
    -196.6200333695437
    -85.40819256078339

     
  2. IMC alignment: Locking nicely. I ran WFS relief to move the WFS output on to the alignment sliders. All the WFS feedback values are now <10. Here is the slider snapshots. [Attachment 2]
     
  3. PMC alignmnet: The PMC looked like it was quite misaligned -> aligned. IMC/PMC locking snapshot [Attachment 3]
    Arm transmissions:
    scripts>z avg 10  C1:LSC-TRX_OUT C1:LSC-TRY_OUT
    C1:LSC-TRX_OUT 0.9825591325759888
    C1:LSC-TRY_OUT 0.9488834202289581

     
  4. Suspension Status Snapshot [Attachment 4]
     
  5. Anchal aligned the OPLEV beams [ELOG 16407]
    I also checked the 100 days trend of the OPLEV sum power. The trend of the max values look flat and fine. [Attachment 5] For this purpose, the PRM and SRM was aligned and the SRM oplev was also aligned. The SRM sum was 23580 when aligned and it was just fine (this is not so visible in the trend plot).
     
  6. The X and Y green beams were aligned for the cavity TEM00s. Y end green PZT values were nulled. The transmission I could reach was as follows.
    >z read C1:ALS-TRX_OUTPUT C1:ALS-TRY_OUTPUT
    0.42343354488901286
    0.24739624058377277

    It seems that these GTRX and GTRY seemed to have crosstalk. When each green shutters were closed the transmissino and the dark offset were measured to be
    >z read C1:ALS-TRX_OUTPUT C1:ALS-TRY_OUTPUT
    0.41822833190834546
    0.025039383697636856
    >z read C1:ALS-TRX_OUTPUT C1:ALS-TRY_OUTPUT
    0.00021112720155274818
    0.2249448773499293

    Note that Y green seemed to have significant (~0.1) of 1st order HOM. I don't know why I could not transfer this power into TEM00. I could not find any significant clipping of the TR beams on the PSL table PDs.
     
  7. IMC Power reduction
    Now we have nice motorized HWP. sitemap -> PSL -> Power control
    == Initial condition == [Attachment 6]
    C1:IOO-HWP_POS 38.83
    Measured input power = 0.99W
    C1:IOO-MC_RFPD_DCMON = 5.38
    == Power reduction == [Attachment 7]
    - The motor was enabled upon rotation on the screen

    C1:IOO-HWP_POS 74.23
    Measured input power = 98mW
    C1:IOO-MC_RFPD_DCMON = 0.537
    - Then, the motor was disabled
     
  8. Went to the detection table and swapped the 10% reflector with the 98% reflector stored on the same table. [Attachments 8/9]
    After the beam alignment the MC REFL PD received about the same amount of the light as before.
    C1:IOO-MC_RFPD_DCMON = 5.6
    There is no beam delivered to the WFS paths.
    CAUTION: IF THE POWER IS INCREASED TO THE NOMINAL WITH THIS CONFIGURATION, MC REFL PD WILL BE DESTROYED.
  9. The IMC can already be locked with this configuration. But for the MC Autolocker, the MCTRANS threshold for the autolocker needs to be reduced as well.
    This was done by swapping a line in  /opt/rtcds/caltech/c1/scripts/MC/AutoLockMC.init
    # BEFORE
    /bin/csh ./AutoLockMC.csh >> $LOGFILE
    #/bin/csh ./AutoLockMC_LowPower.csh >> $LOGFILE
    --->
    # AFTER
    #/bin/csh ./AutoLockMC.csh >> $LOGFILE
    /bin/csh ./AutoLockMC_LowPower.csh >> $LOGFILE

    Confirmed that the autolocker works a few times by toggling the PSL shutter. The PSL shutter was closed upon the completion of the test
     
  10. Walked around the lab and checked all the bellows - the jam nuts are all tight, and I couldn't move them with my hands. So this is okay according to the ancient tale by Steve.
Attachment 1: Screenshot_2021-10-15_17-36-00.png  27 kB  Uploaded Fri Oct 15 18:36:33 2021  | Hide | Hide all
Screenshot_2021-10-15_17-36-00.png
Attachment 2: Screenshot_2021-10-15_17-39-58.png  13 kB  Uploaded Fri Oct 15 18:40:22 2021  | Hide | Hide all
Screenshot_2021-10-15_17-39-58.png
Attachment 3: Screenshot_2021-10-15_17-42-20.png  27 kB  Uploaded Fri Oct 15 18:43:06 2021  | Hide | Hide all
Screenshot_2021-10-15_17-42-20.png
Attachment 4: Screenshot_2021-10-15_17-46-13.png  26 kB  Uploaded Fri Oct 15 18:46:21 2021  | Hide | Hide all
Screenshot_2021-10-15_17-46-13.png
Attachment 5: Screenshot_2021-10-15_18-05-54.png  49 kB  Uploaded Fri Oct 15 19:06:48 2021  | Hide | Hide all
Screenshot_2021-10-15_18-05-54.png
Attachment 6: Screen_Shot_2021-10-15_at_19.45.05.png  121 kB  Uploaded Fri Oct 15 21:43:32 2021  | Hide | Hide all
Screen_Shot_2021-10-15_at_19.45.05.png
Attachment 7: Screen_Shot_2021-10-15_at_19.47.10.png  135 kB  Uploaded Fri Oct 15 21:43:42 2021  | Hide | Hide all
Screen_Shot_2021-10-15_at_19.47.10.png
ELOG V3.1.3-